OpenStudy (anonymous):

Which of the following represents the area of a rectangle whose length is x + 1 and whose width is x + 11? x^2+ 12x + 11 x^2+ 11 x^2+ 11x + 12 x^2+ 10x + 11 Which of the following represents the area of a rectangle whose length is x + 1 and whose width is x + 11? x^2+ 12x + 11 x^2+ 11 x^2+ 11x + 12 x^2+ 10x + 11 @Mathematics

OpenStudy (hoblos):

x^2+ 12x + 11

OpenStudy (anonymous):

Area of rectangle = length * breadth/width = (x + 1) * (x + 11) = x(x + 11) +1(x + 11) = x^2 + 11x + x + 11 = x^2 + 12x + 11